OpenStudy (anonymous):

Use elimination to solve the system of equations. Enter your answer as an ordered pair; example: (4, -5) 1/2x+y=4 -2x-2y=-6

OpenStudy (mertsj):

Multiply the first equation by 2 x+2y=8 Add the second equation: -x=2 x=-2 Substitute into the first equation: 1/2(-2)+y=4 -1+y=4 y=5